OpenStudy (anonymous):

what does cot(pi^2/16) equal?

OpenStudy (badhi):

$$\begin{align*} \cot \left(\frac{\pi^2}{16}\right)&=\frac{\cos\left(\frac{\pi^2}{16}\right)}{\sin\left(\frac{\pi^2}{16}\right)}\\ &=\frac{\cos \left(2\frac{\pi}{4}\right)+1}{1-\cos \left(2\frac{\pi}{4}\right)}\\ &=\frac{\cos\left(\frac \pi 2\right)+1}{1-\cos\left(\frac \pi 2\right)} \end{align*}$$

OpenStudy (anonymous):

Awesome! So the original question was to get the derivative at a, (a=-pi/4) for the function \[\ f(x)=sin^{-1}(x ^{2})-1/\sin^2x+2\sin(2x) \] I guess I'm looking for a sort of exact answer?